Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $40 per unit. Variable expenses are $20.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

Sales $ 1,040,000
Variable expenses 520,000
Contribution margin 520,000
Fixed expenses 180,000
Net operating income $ 340,000

The sales manager is convinced that a 11% reduction in the selling price, combined with a $74,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?
Answer #1

Solution a:

New selling price per unit = $40 * 89% = $35.60

New Fixed costs = $180,000 + $74,000 = $254,000

New sales volume = ($1,040,000 / $40)*125% = 32500 units

Computation of net operating income - Feather friend Inc.
Particulars Amount
Sales (32500 * $35.6) 1157000
Variable expenses (32500*$20) 650000
Contribution margin 507000
Fixed expenses 254000
Net operating income 253000

Solution b:

net operating income decrease by = $253000 - 340000 = ($87,000)